# LeanRAG: Knowledge-Graph-Based Generation with Semantic Aggregation and Hierarchical Retrieval

### v2

@@ Abstract
Retrieval-Augmented Generation (RAG) plays a crucial role in grounding Large Language Models by leveraging external knowledge, whereas the effectiveness is often compromised by the retrieval of contextually flawed or incomplete information. To address this, knowledge graph-based RAG methods have evolved towards hierarchical structures, organizing knowledge into multi-level summaries. However, these approaches still suffer from two critical, unaddressed challenges: high-level conceptual summaries exist as disconnected “semantic islands”, lacking the explicit relations needed for cross-community reasoning; and the retrieval process itself remains structurally unaware, often degenerating into an inefficient flat search that fails to exploit the graph’s rich topology. To overcome these limitations, we introduce LeanRAG, a framework that features a deeply collaborative design combining knowledge aggregation and retrieval strategies. LeanRAG first employs a novel semantic aggregation algorithm that forms entity clusters and constructs new explicit relations among aggregation-level summaries, creating a fully navigable semantic network. Then, a bottom-up, structure-guided retrieval strategy anchors queries to the most relevant fine-grained entities and then systematically traverses the graph’s semantic pathways to gather concise yet contextually comprehensive evidence sets. The LeanRAG can mitigate the substantial overhead associated with path retrieval on graphs and minimizes redundant information retrieval. Extensive experiments on four challenging QA benchmarks with different domains demonstrate that LeanRAG significantly outperforming existing methods in response quality while reducing 46% retrieval redundancy. Code is available at: https://github.com/RaZzzyz/LeanRAG
@@ Introduction
To overcome the limitations of unstructured retrieval, researchers have increasingly explored knowledge graph based RAG methods. Initial efforts, such as GraphRAG (Edge et al. 2024), successfully organized documents into community-based knowledge graphs, which helped preserve local context better than disconnected text chunks. However, these methods often generated large, coarse-grained communities, leading to significant information redundancy during retrieval. Subsequent, more advanced works like HiRAG (Huang et al. 2025a) refined this paradigm by introducing hierarchical structures, clustering entities into multi-level summaries. This represented a significant step forward in organizing knowledge. Despite this progress, our analysis reveals that two critical challenges remain unaddressed currently (as Figure 1 shows). First, the high-level summary nodes in these hierarchies exist as “semantic islands”. They lack explicit relational connections between each other, making it hard to reason across different conceptual communities within the knowledge base. Second, the retrieval process itself remains structurally unaware, often degenerating into a simple semantic search over a flattened list of nodes, failing to exploit the rich topological information encoded in the graph. This leads to a retrieval process that is both inefficient and imprecise.
@@
Building upon this enriched structure, LeanRAG employs a bottom-up, structure-aware retrieval mechanism that strategically navigates the graph to maximize relevance while minimizing redundancy. The retrieval process begins by anchoring the query to the most contextually pertinent fine-grained entities at the leaf level. It then systematically traverses relational pathways across both the original entity layer and the derived summary layer, propagating evidence upward through the hierarchy. This dual-level traversal ensures that the retrieved evidence set is not only concise and focused, but also contextually comprehensive, capturing both specific details and broader conceptual relations essential for accurate and coherent generation.
@@
  • The introduction of a bottom-up entities retrieval strategy to mitigate information redundancy. By initiating retrieval from high-relevance “anchor” nodes and expanding context strictly along relevant semantic pathways, this strategy yields a precise and compact evidence subgraph for LLMs.
@@
Substantial research has been dedicated to overcoming this limitation. One line of work focuses on improving the retriever itself, evolving from sparse methods like BM25 (Robertson, Zaragoza et al. 2009) to dense retrieval models such as DPR (Karpukhin et al. 2020) and Contriever (Izacard et al. 2021), which learn to better capture semantic relevance. Another direction targets the indexing and organization of the source documents (Jiang et al. 2023). Recent advancements have explored creating hierarchical summaries of text chunks, allowing retrieval to occur at multiple levels of granularity. For instance, RAPTOR builds a tree of recursively summarized text clusters, enabling retrieval of both fine-grained details and high-level summaries (Sarthi et al. 2024). Despite these improvements, these methods still largely treat knowledge as a linear sequence or a simple tree of text. They do not explicitly model the complex, non-hierarchical relations that often exist between different entities and concepts within a document. This limitation hinders their ability to answer complex queries that require reasoning over these intricate connections, motivating the shift towards KG-based RAG methods.
@@
In this section, we will introduce and give a formal definition of a RAG system with specific knowledge graph.
Given a rich knowledge graph with the description of vertexes and relations $\mathcal{G}=(V,R,D_{\text{(}ver)},D_{\text{(}rel)})$, where $V$ and $R$ denote the set of entities and relations, $D_{\text{(}ver)}$ represents the collection of entity descriptions and $D_{\text{(}rel)}$ represents the collection of relationship descriptions. The goal of KG-based RAG is to leverage existing information to build a query-relevant sub-graph which helps LLMs generate high-quality response. Given a query $q$, the searching process can be formulated as:
@@
where $Sim(\cdot,\cdot)$ is the embedding similarity metric function, and n is the choice number of similarity entity. Based on the metric, $\tilde{V}$ contains the top n entities. Then we can search the relational paths $L$ between nodes $v\in\tilde{V}$. All relations $r$ that constitute the path $L$ belong to the relation set $R$.
@@
The performance of a generic KG-augmented retrieval framework is fundamentally determined by the structural and semantic quality of underlying knowledge graph $\mathcal{G}$, as well as the precision and efficiency of the retrieval strategy. To address the limitations of a flat graph structure and naive path search strategy, we introduce LeanRAG, a framework built on the principle of tightly co-designing its aggregation and retrieval processes. As illustrated in Figure 2, LeanRAG consists of two core innovations: (1) a Hierarchical Graph Aggregation method that recursively builds a multi-level, navigable semantic network from the base graph; and (2) a Structured Retrieval strategy that leverages this hierarchy via Lowest Common Ancestor (LCA) path search approach to construct a compact and coherent context.
@@
Given a knowledge graph layer $\mathcal{G}_{i-1}$, the first step is to identify groups of semantically related entities that can be abstracted into a single, higher-level concept. We leverage the rich descriptive text $d_{v}\in D_{{\text{(}ver)}_{i-1}}$ associated with each entity $v\in V_{i-1}$ for this purpose. Following recent successes in clustering text representation (Sarthi et al. 2024), we employ a two-step process:
@@
In practice, the generation function $\mathcal{F}_{\text{entity}}$ is implemented by LLMs guided by a carefully designed prompt $\mathcal{P}_{\text{entity}}$, details are shown in Appendix D.1. We prompt LLMs to produce a concise name for the new entity $\alpha_{j}$ and a comprehensive description $d_{\alpha_{j}}$ that summarizes its components. Each entity $v\in C_{j}$ is then linked to its new parent entity $\alpha_{j}$, forming the parent-child connections in the hierarchy.
@@
In practice, the generation function $\mathcal{F}_{\text{rel}}$ is implemented by LLMs guided by a specific prompt $\mathcal{P}_{\text{rel}}$, details are shown in Appendix D.2.
@@
Given a user query $q$, the first step is to ground the query in the most specific, fine-grained facts available. We achieve this by performing a dense retrieval search exclusively over the entities of the original graph including the initial entities, that is, the base-layer graph $\mathcal{G}_{0}$. We identify the top n entities whose textual descriptions are most semantically similar to the query:
@@
This set of “seed entities”, $V_{\text{seed}}$, serves as the starting point for structured traversal, ensuring our retrieval process is anchored in the most relevant parts of knowledge base.
@@
where $R_{\text{lca}}$ contains the relations within the retrieval path $\mathcal{P}_{\text{lca}}$ and $R_{\text{inter-cluster}}$ contains the inter-cluster relations between aggregation entities which are in the same level in the hierarchical knowledge graph. For example, $r_{<\alpha_{j},\alpha_{k}>}\in R_{\text{inter-cluster}}$, where $\alpha_{j}\in\mathcal{G}_{i}$ and $\alpha_{k}\in\mathcal{G}_{i}$.
@@
  • RQ4: Is the structured knowledge retrieved from the graph sufficient for high-quality generation, or is the inclusion of the entities’ original textual context essential?
@@
  • LightRAG (Guo et al. 2024): Employs a dual-level retrieval framework built upon a KG-based text indexing paradigm.
@@
To provide a multi-faceted and in-depth analysis of system performance, we evaluate the generated answers along four crucial dimensions follows the prior work (Huang et al. 2025a):
@@
  • Overall: Provides a single, holistic quality score to measure how the answer perform overall, considering comprehensiveness, empowerment, diversity, and any other relevant factors.
Following recent best practices in automated evaluation, we employ powerful LLMs as judges to score the outputs of all methods on the 1 to 10 scale defined by our metrics. In order to directly reflect the quality of the answers, we will also use LLM to directly evaluate the two answers to obtain their win rates. Specifically, we use DeepSeek-V3 (Liu et al. 2024) as our evaluators, providing them with carefully designed prompts to ensure consistent and unbiased scoring, and each query and answer is scored at 5 times. Detail prompt is shown in Appendix D.3.
@@
Across all experiments, we use DeepSeek-V3 as LLM generator for all models to ensure a fair comparison. The text embedding for retrieval is computed using BGE-M3 (Chen et al. 2024). The number of clusters for the GMM and other key hyperparameters are tuned on a held-out validation set. All main experiments were conducted by leveraging commercial API services. For our main experiments, we utilized the Deepseek-V3 model as the backbone for all models follow the prior work (Huang et al. 2025a), ensuring a fair comparison. In addition, in order to evaluate RQ2 efficiently, we reproduced the baseline methods on the Qwen3-14b (Yang et al. 2025) model to evaluate the redundancy between LeanRAG and other methods. Details are shown in Appendix B.
@@
To address RQ1 , we compare LeanRAG against all baseline models across four benchmarks, as presented in Table 1. The experimental results demonstrate that LeanRAG almost outperforms all baselines across the evaluated datasets. And we provide the results of winrate metric in the Appendix A.
Table 1: Evaluation scores (1–10 scale) of LeanRAG compared to baseline methods, assessed by a LLM
@@
From a Comprehensiveness perspective, even after removing the information-intensive community structure of traditional KG-based RAG, the aggregation used by LeanRAG still provides sufficient query-related information. Furthermore, Empowerment and Diversity effectively measure the relevance of the provided information. These indicate that LeanRAG effectively enhances the breadth of information by establishing inter-cluster relations, resulting in optimal performance. In summary, LeanRAG demonstrates state-of-the-art performance on the majority of metrics across four evaluated datasets, and achieves highly competitive results on the remaining ones.
Analysis of Information Rebundancy (RQ2)
@@
Table 2: Win rates (%) between LeanRAG and LeanRAG wo/relation (Left: LeanRAG; Right: w/o Relation)
@@
To answer RQ4, we investigate the role of the original, unstructured text chunks in our framework. While our graph structure serves as an effective retrieval guide, it is crucial to understand whether the structured information alone is sufficient for the generator, or if the source text is essential. To this end, we conduct an ablation study by creating a variant of our model, denoted as LeanRAG w/o Context. This variant performs the exact same hierarchical retrieval process, but the final context provided to the LLM generator consists only of the names and descriptions of the retrieved graph entities, excluding the original text chunks associated with the base-level entities. We then compare its performance against the full LeanRAG model.
@@
The most pronounced drops are consistently seen in the Comprehensiveness and Empowerment metrics. This is expected, as the raw text chunks contain the detailed explanations, evidence, and nuanced language necessary for generating thorough and actionable answers. In contrast, a context composed solely of structured entity information, while semantically focused, lacks the narrative richness required by the LLM. These findings confirm our hypothesis: the hierarchical graph in LeanRAG acts as a highly effective semantic index and navigation system whose primary function is to precisely locate the most critical segments of unstructured text. The collaboration between the structured graph traversal for guidance and the rich content of the unstructured text for generation is essential to achieving state-of-the-art performance.
Table 3: Necessity analysis of textual context
@@ Conclusions
To addressed the critical challenges of “semantic islands” and the structure-retrieval mismatch prevalent in the KG-based RAG systems, we introduced LeanRAG, a novel framework that resolves these issues through a tight co-design of its knowledge aggregation and retrieval mechanisms. Our approach features a hierarchical aggregation algorithm that constructs a fully navigable semantic network by generating explicit relations between abstract summary concepts, and a complementary bottom-up, LCA-based retrieval strategy that efficiently traverses this structure. Extensive experiments validated our design, demonstrating that LeanRAG achieves state-of-the-art performance while significantly reducing information redundancy. Furthermore, our ablation studies confirmed that both the generation of summary information and original textual context are essential for producing comprehensive and diverse answers.

### v4

@@ Abstract
Retrieval-Augmented Generation (RAG) plays a crucial role in grounding Large Language Models by leveraging external knowledge, whereas the effectiveness is often compromised by the retrieval of contextually flawed or incomplete information. To address this, knowledge graph-based RAG methods have evolved towards hierarchical structures, organizing knowledge into multi-level summaries. However, these approaches still suffer from two critical, unaddressed challenges: high-level conceptual summaries exist as disconnected “semantic islands”, lacking the explicit relations needed for cross-community reasoning; and the retrieval process itself remains structurally unaware, often degenerating into an inefficient flat search that fails to exploit the graph’s rich topology. To overcome these limitations, we introduce LeanRAG, a framework that features a deeply collaborative design combining knowledge aggregation and retrieval strategies. LeanRAG first employs a novel semantic aggregation algorithm that forms entity clusters and constructs new explicit relations among aggregation-level summaries, creating a fully navigable semantic network. Then, a bottom-up, structure-guided retrieval strategy anchors queries to the most relevant fine-grained entities and then systematically traverses the graph’s semantic pathways to gather concise yet contextually comprehensive evidence sets. The LeanRAG can mitigate the substantial overhead associated with path retrieval on graphs and minimize redundant information retrieval. Extensive experiments on four challenging QA benchmarks with different domains demonstrate that LeanRAG significantly outperforms existing methods in response quality while reducing 46% retrieval redundancy. Our code is available at: https://github.com/RaZzzyz/LeanRAG.
@@ Introduction
To overcome the limitations of unstructured retrieval, researchers have increasingly explored knowledge graph-based RAG methods. Initial efforts, such as GraphRAG (Edge et al. 2024), successfully organized documents into community-based knowledge graphs, which helped preserve local context better than disconnected text chunks. However, these methods often generated large, coarse-grained communities, leading to significant information redundancy during retrieval. Subsequently, more advanced works like HiRAG (Huang et al. 2025a) refined this paradigm by introducing hierarchical structures, clustering entities into multi-level summaries. This represented a significant step forward in organizing knowledge. Despite this progress, our analysis reveals that two critical challenges remain unaddressed currently (as Figure 1 shows). First, the high-level summary nodes in these hierarchies exist as “semantic islands”. They lack explicit relational connections between each other, making it hard to reason across different conceptual communities within the knowledge base. Second, the retrieval process itself remains structurally unaware, often degenerating into a simple semantic search over a flattened list of nodes, failing to exploit the rich topological information encoded in the graph. This leads to a retrieval process that is both inefficient and imprecise.
@@
Building upon this enriched structure, LeanRAG employs a bottom-up, structure-aware retrieval mechanism that strategically navigates the graph to maximize relevance while minimizing redundancy. The retrieval process begins by anchoring the query to the most contextually pertinent fine-grained entities at the leaf level. It then systematically traverses relational pathways across both the original entity layer and the derived summary layer, propagating evidence upward through the hierarchy. This dual-level traversal ensures that the retrieved evidence set is not only concise and focused but also contextually comprehensive, capturing both specific details and broader conceptual relations essential for accurate and coherent generation.
@@
  • The introduction of a bottom-up entity retrieval strategy to mitigate information redundancy. By initiating retrieval from high-relevance “anchor” nodes and expanding context strictly along relevant semantic pathways, this strategy yields a precise and compact evidence subgraph for LLMs.
@@
Substantial research has been dedicated to overcoming this limitation. One line of work improves the retriever itself, evolving from sparse methods like BM25 (Robertson, Zaragoza et al. 2009) to dense models such as DPR (Karpukhin et al. 2020) and Contriever (Izacard et al. 2021), which better capture semantic relevance. Another focuses on indexing and organizing source documents (Jiang et al. 2023), with recent methods creating hierarchical summaries of text chunks to enable multi-level retrieval. For example, RAPTOR builds a tree of recursively summarized clusters, allowing retrieval of fine-grained details and high-level summaries (Sarthi et al. 2024). However, these approaches still treat knowledge as linear or simple hierarchical structures and do not explicitly model complex, non-hierarchical relations between entities and concepts, limiting their ability to answer queries requiring reasoning over such connections—motivating KG-based RAG methods.
@@
In this section, we will introduce and give a formal definition of a RAG system with a specific knowledge graph.
Given a rich knowledge graph with the description of vertexes and relations $\mathcal{G}=(V,R,D_{\text{(}ver)},D_{\text{(}rel)})$, where $V$ and $R$ denote the set of entities and relations, $D_{\text{(}ver)}$ represents the collection of entity descriptions and $D_{\text{(}rel)}$ represents the collection of relationship descriptions. The goal of KG-based RAG is to leverage existing information to build a query-relevant sub-graph that helps LLMs generate high-quality responses. Given a query $q$, the searching process can be formulated as:
@@
where $Sim(\cdot,\cdot)$ is the embedding similarity metric function, and n is the choice number of similarity entities. Based on the metric, $\tilde{V}$ contains the top n entities. Then we can search the relational paths $L$ between nodes $v\in\tilde{V}$. All relations $r$ that constitute the path $L$ belong to the relation set $R$.
@@
The performance of a generic KG-augmented retrieval framework is fundamentally determined by the structural and semantic quality of the underlying knowledge graph $\mathcal{G}$, as well as the precision and efficiency of the retrieval strategy. To address the limitations of a flat graph structure and naive path search strategy, we introduce LeanRAG, a framework built on the principle of tightly co-designing its aggregation and retrieval processes. As illustrated in Figure 2, LeanRAG consists of two core innovations: (1) a Hierarchical Graph Aggregation method that recursively builds a multi-level, navigable semantic network from the base KG; and (2) a Structured Retrieval strategy that leverages this hierarchy via Lowest Common Ancestor (LCA) path search approach to construct a compact and coherent context.
@@
Given a knowledge graph layer $\mathcal{G}_{i-1}$, the first step is to identify groups of semantically related entities that can be abstracted into a single, higher-level concept. We leverage the rich descriptive text $d_{v}\in D_{{\text{(}ver)}_{i-1}}$ associated with each entity $v\in V_{i-1}$ for this purpose. Following recent works in clustering text representation (Sarthi et al. 2024), we employ a two-step process:
@@
In practice, the generation function $\mathcal{F}_{\text{entity}}$ is implemented by LLMs guided by a carefully designed prompt $\mathcal{P}_{\text{entity}}$. We prompt LLMs to produce a concise name for the new entity $\alpha_{j}$ and a comprehensive description $d_{\alpha_{j}}$ that summarizes its components. Each entity $v\in C_{j}$ is then linked to its new parent entity $\alpha_{j}$, forming the parent-child connections in the hierarchy.
@@
In practice, the generation function $\mathcal{F}_{\text{rel}}$ is implemented by LLMs guided by a specific prompt $\mathcal{P}_{\text{rel}}$.
@@
Given a user query $q$, the first step is to ground the query in the most specific, fine-grained facts available. We achieve this by performing a dense retrieval search exclusively over the entities of the original graph, including the initial entities, that is, the base-layer graph $\mathcal{G}_{0}$. We identify the top n entities whose textual descriptions are most semantically similar to the query:
@@
This set of “seed entities”, $V_{\text{seed}}$, serves as the starting point for structured traversal, ensuring our retrieval process is anchored in the most relevant parts of the knowledge base.
@@
where $R_{\text{lca}}$ contains the relations within the retrieval path $\mathcal{P}_{\text{lca}}$ and $R_{\text{inter-cluster}}$ contains the inter-cluster relations between aggregation entities that are in the same level in the hierarchical knowledge graph. For example, $r_{<\alpha_{j},\alpha_{k}>}\in R_{\text{inter-cluster}}$, where $\alpha_{j}\in\mathcal{G}_{i}$ and $\alpha_{k}\in\mathcal{G}_{i}$.
@@
  • RQ4: Is the structured knowledge retrieved from the graph sufficient for high-quality generation, or is the inclusion of the entities original textual context essential?
@@
  • LightRAG (Guo et al. 2024): Uses a dual-level retrieval framework based on a KG-based text indexing paradigm.
@@
To provide a multi-faceted and in-depth analysis of system performance, we evaluate the generated answers along four crucial dimensions, following the prior work (Huang et al. 2025a):
@@
  • Overall: Provides a single, holistic quality score to measure how the answer performs overall, considering comprehensiveness, empowerment, diversity, and any other relevant factors.
Following recent best practices in automated evaluation, we employ powerful LLMs as judges to score the outputs of all methods on the 1 to 10 scale defined by our metrics. In order to directly reflect the quality of the answers, we will also use LLM to directly evaluate the two answers to obtain their win rates. Specifically, we use DeepSeek-V3 (Liu et al. 2024) as our evaluators, providing them with carefully designed prompts to ensure consistent and unbiased scoring, and each query and answer is scored 5 times.
@@
Across all experiments, we use DeepSeek-V3 as an LLM generator for all models to ensure a fair comparison. The text embedding for retrieval is computed using BGE-M3 (Chen et al. 2024). The number of clusters for the GMM and other key hyperparameters are tuned on a held-out validation set. All main experiments were conducted by leveraging commercial API services. For our main experiments, we utilized the Deepseek-V3 model as the backbone for all models, following the prior work (Huang et al. 2025a), ensuring a fair comparison. In addition, in order to evaluate RQ2 efficiently, we reproduced the baseline methods on the Qwen3-14b (Yang et al. 2025) model to evaluate the redundancy between LeanRAG and other methods. All implementation details are provided in the extended version (zhang2025leanragknowledgegraphbasedgenerationsemantic).
@@
To address RQ1 , we compare LeanRAG against all baseline models across four benchmarks, as presented in Table 1. The experimental results demonstrate that LeanRAG almost outperforms all baselines across the evaluated datasets.
@@
Table 1: Evaluation scores (1–10 scale) of LeanRAG compared to baseline methods, assessed by an LLM
From a Comprehensiveness perspective, even after removing the information-intensive community structure of traditional KG-based RAG, the aggregation used by LeanRAG still provides sufficient query-related information. Furthermore, Empowerment and Diversity effectively measure the relevance of the provided information. These indicate that LeanRAG effectively enhances the breadth of information by establishing inter-cluster relations, resulting in optimal performance. In summary, LeanRAG demonstrates state-of-the-art performance on the majority of metrics across four evaluated datasets and achieves highly competitive results on the remaining ones.
Analysis of Information Redundancy (RQ2)
@@
Table 2: Win rates (%) between LeanRAG and LeanRAG w/o Relation (Left: LeanRAG; Right: w/o Relation)
@@
To answer RQ4, we investigate the role of the original unstructured text chunks in our framework. While the graph structure serves as an effective retrieval guide, it is crucial to assess whether structured information alone suffices for the generator or if the source text remains essential. To this end, we conduct an ablation study using a variant of our model, denoted as LeanRAG w/o Context. This variant follows the same hierarchical retrieval process, but the final context provided to the LLM generator includes only the names and descriptions of the retrieved graph entities, excluding the original text chunks linked to the base-level entities. We then compare its performance with that of the full LeanRAG model.
@@
The most pronounced drops are consistently seen in the Comprehensiveness and Empowerment metrics. This is expected, as raw text chunks contain the detailed explanations, evidence, and nuanced language necessary for generating thorough and actionable answers. In contrast, a context composed solely of structured entity information, while semantically focused, lacks the narrative richness required by the LLM. These findings confirm our hypothesis: the hierarchical graph in LeanRAG acts as an effective semantic index and navigation system whose primary function is to precisely locate critical segments of unstructured text. The collaboration between structured graph traversal for guidance and the rich content of unstructured text for generation is essential to achieving state-of-the-art performance.
@@ Conclusions
To address the critical challenges of “semantic islands” and the structure-retrieval mismatch prevalent in the KG-based RAG systems, we introduced LeanRAG, a novel framework that resolves these issues through a tight co-design of its knowledge aggregation and retrieval mechanisms. Our approach features a hierarchical aggregation algorithm that constructs a fully navigable semantic network by generating explicit relations between abstract summary concepts, and a complementary bottom-up, LCA-based retrieval strategy that efficiently traverses this structure. Extensive experiments validated our design, demonstrating that LeanRAG achieves state-of-the-art performance while significantly reducing information redundancy. Furthermore, our ablation studies confirmed that both the generation of summary information and the original textual context are essential for producing comprehensive and diverse answers.
Acknowledgments
The research was supported by Shanghai Artificial Intelligence Laboratory, the National Key R&D Program of China (Grant No. 2022ZD0160201) and the Science and Technology Commission of Shanghai Municipality (Grant Nos. 22DZ1100102).